OpenStudy (anonymous):

Reggie is three less than four times his daughter's age. Reggie is 69; how old is his daughter? a.) 10 B.) 18 C.) 12 D.) 20 help?

OpenStudy (kropot72):

Let d be the daughter's age. Then 69 = 4d - 3 ...........(1) Just solve equation (1) to find d, the daughter's age.

OpenStudy (anonymous):

Reggie = 69 (R = 69) Reggis is 3 less (R = -3) than four times his Daughter's age (R = 4D - 3) So, 69 = 4D - 3

OpenStudy (anonymous):

69 + 3 = 4D 72 = 4D 72/4 = D